Calculate the following.

    

a. The future value of $460 eight years from now at 7 percent. (Round your final answer to 2 decimal places.)

b. The future value of $500 saved each year for 7 years at 5 percent. (Round your final answer to 2 decimal places.)

  

c. The amount a person would have to deposit today (present value) at an interest rate of 6 percent to have $2,200 five years from now. (Round your final answer to 2 decimal places.)

  

d. The amount a person would have to deposit today to be able to take out $500 a year for 8 years from an account earning 5 percent. (Round your final answer to 2 decimal places.)
Answer #1

We know that,

Future Value = Present Value * (1+ interest rate)^ years

a.)

Future Value = Present Value * (1+ interest rate)^ years

Future Value = 460 * (1+ 0.07)^8 = 790.37 Answer

b.)

Total Future Value = Sum of future value of all the payments

Future Value = 500 *(1+0.05)^(7-1) +500 *(1+0.05)^(7-2) +500 *(1+0.05)^(7-3) +500 *(1+0.05)^(7-4) +500 *(1+0.05)^(7-5) +500 *(1+0.05)^(7-6) +500 *(1+0.05)^(7-7)

Future Value = 4071 Answer

c.)

Future Value = Present Value * (1+ interest rate)^ years

2200 = Present Value * (1+0.06)^5

Present Value = 1643.97 Answer

d.)

Amount a person would have to deposit = 500/(1+0.05)^1 +500/(1+0.05)^2 +500/(1+0.05)^3 +500/(1+0.05)^4 +500/(1+0.05)^5 +500/(1+0.05)^6 +500/(1+0.05)^7 +500/(1+0.05)^8

Amount a person would have to deposit = 3231.61 Answer

Kindly do inform me in case you have any queries.